OpenStudy (anonymous):

open cover of a set is denoted by \(\left\{ U_{\alpha} \right\}\). What does alpha stand for?

OpenStudy (zzr0ck3r):

maybe an index?

ganeshie8 (ganeshie8):

indexing set i think @zzr0ck3r

OpenStudy (anonymous):

can you give examples? like what does \(\left\{ U_{\pi} \right\}\)

OpenStudy (anonymous):

...stand for?

OpenStudy (zzr0ck3r):

um its just a stand in for an index so the amount of sets you are taking the union of is \(\alpha\)

OpenStudy (zzr0ck3r):

\(\pi\) would not make sense if we are talking about 3.14....

OpenStudy (zzr0ck3r):

but we use pi for many things...

OpenStudy (anonymous):

but alpha doesn't have to be a natural number tho?

OpenStudy (zzr0ck3r):

no

OpenStudy (zzr0ck3r):

you can have infinite and uncountable infinite

OpenStudy (zzr0ck3r):

http://en.wikipedia.org/wiki/Cover_(topology)

OpenStudy (zzr0ck3r):

Just really bad short notation....

ganeshie8 (ganeshie8):

if i remember correctly you could define alpha as an interval \[\alpha_n = (1, n)\] then the set \(\{U_{\alpha_{1}}\}\) refers to the set \((1, 1)\)

OpenStudy (anonymous):

let's say the interval [0,1]. How would you denote that for every element in [0,1] take a ball of radius 1/10?

ganeshie8 (ganeshie8):

@zzr0ck3r please correct me if i talk nonsense :)

OpenStudy (zzr0ck3r):

Yeah that seems fine to me. its just notation...

ganeshie8 (ganeshie8):

should be like this i think then the set \(\{U_{\alpha_{1}}\}\) refers to the set \(\{(1, 1)\}\)

OpenStudy (zzr0ck3r):

so you want the collection of balls around x s.t. x is in that set?

OpenStudy (anonymous):

yes @zzr0ck3r

OpenStudy (anonymous):

so is alpha is like the notation for sequence a(n) where a itself does not represent a number?

OpenStudy (zzr0ck3r):

alpha is really bad notation for this \(\cup_{i\in \alpha}A_i\)

OpenStudy (zzr0ck3r):

where A_i is one of the setes in the collection and there are \(\alpha\) sets

OpenStudy (zzr0ck3r):

\(\{B(x,\frac{1}{10}) \mid x\in [0,1]\}\)

OpenStudy (anonymous):

:O so alpha is a set?

OpenStudy (zzr0ck3r):

yes, an indexed set

OpenStudy (watchmath):

\(\alpha\) is a member of an index set, say \(\Gamma\)

OpenStudy (zzr0ck3r):

\(\alpha\) is the indexed set

OpenStudy (zzr0ck3r):

we would never just list a member of an index as subscript in a union, there would be no point

OpenStudy (anonymous):

the whole time i thought it was a number! -.- that's why it got even more confused when it comes to the notation for subcover

OpenStudy (zzr0ck3r):

every number is a set ;)

OpenStudy (watchmath):

I mean in your example of your open cover \(\{U_\alpha\}\), \(\alpha\) is a member of some index set.

OpenStudy (zzr0ck3r):

no, I think alpha is the indexed set, why would you union the elements of a set, you just get back the set.....

OpenStudy (zzr0ck3r):

http://en.wikipedia.org/wiki/Cover_(topology)

OpenStudy (zzr0ck3r):

second line

OpenStudy (zzr0ck3r):

its bad notation lol

OpenStudy (watchmath):

wikipedia is right, \(\alpha\) is a member of index set \(A\) there

OpenStudy (zzr0ck3r):

but this short notation stands for the union of the indexed set.

OpenStudy (zzr0ck3r):

read the next line

OpenStudy (watchmath):

the next line is consistent, the collection of \(U_\alpha\) meaning \(\alpha\) runs through some index set. If you have learned topology before, it is a common notation that \(\alpha\) is a member of an index set.

OpenStudy (watchmath):

For example you can write \(\{U_\alpha\mid \alpha \in [0,1]\}\) or \(\bigcup_{\alpha \in [0,1]}U_\alpha\)

OpenStudy (zzr0ck3r):

I would never write it like this and this is why I say its bad notation. I would write it as \(\cup_{I\in I}A_i\)

OpenStudy (zzr0ck3r):

err \(\cup_{i\in I}A_i \)

OpenStudy (zzr0ck3r):

they wrote \(\cup_i\)

OpenStudy (watchmath):

well, there is no different with your notation. They just don't specify, what is the exact index set, but that is ok, to save space and time.

OpenStudy (zzr0ck3r):

I am saying what if we have sets \(A_1=\{1,2,3,4\},A_2=\{a,b,c,d\}\) and lets say our index is then explin to me wht \(U_1\) is?

OpenStudy (zzr0ck3r):

where our index is \(I=\{1,2\}\)

OpenStudy (zzr0ck3r):

They don't specify the collection either...

OpenStudy (zzr0ck3r):

so again this is just bad notation for what I said. It does not any longer stand for an element or we would just have \(\cup_1=\cup A_1=A_1\)

OpenStudy (watchmath):

Yes, by you give a name to your set of index and one can avoid it either

OpenStudy (watchmath):

Let say by saying \(\bigcup_\alpha A_\alpha\)

OpenStudy (watchmath):

In this notation it is understood that you have an index set in the background where the \(\alpha\) is a particular member in this set.